Circular No. 9011

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

η CARINAE

A. Damineli, Instituto de Astronomico e Geofísico, Universidade de São Paulo, reports that spectroscopic observations collected at the Southern Astrophysical Research Telescope (La Serena, Chile) by Damineli, J. Steiner, J. H. Groh, and L. Fraga indicate that η Car reached the phase of minimum excitation on Jan. 11, on schedule with the predictions (Damineli et al. 2008, MNRAS 384, 1649). This supports the 2023-day period and the binary nature of the system. The doubly ionized forbidden lines and the narrow components of He I and [N II] lines disappeared from the spectrum. P-Cyg absorption in permitted transitions (He I, H I, and Fe II) became more enhanced. The He II 468.6-nm line, which had been increasing rapidly through 2008 December, has now collapsed abruptly to ∼ 300 mÅ, the same level as it was a month before. The radial velocity reached −3204 km/s and seems to be still shifting toward the blue. Such results are corroborated by observations taken at Observatório do Picos dos Dias (Brazil) by M. Teodoro; at the Complejo Astronómico el Leoncito (Argentina) by E. F. Lajuz, F. Gonzalez, and R. Gamen; at the European Southern Observatory 2.2-m telescope by M. Borges Fernandes, C. B. Pereira, C. A. P. C. O. Torres, and M. I. Zevallos Herencia; and at Las Campanas Observatory by N. Morrell. Figures and additional details can be found at the following

website URL: www.etacarinae.iag.usp.br.

COMETS C/2008 R8–R12 (SOHO) Further to IAUC 9006, additional near-sun presumed comets have been

found on SOHO website images — all Kreutz sungrazers except for C/2008 R8 (non-group). C/2008 R8 (peak mag ∼ 7), C/2008 R9 (mag ∼ 7), and 2008 R11 (mag ∼ 6.5) were small and stellar in appearance. C/2008 R10 was elongated (mag ∼ 5.5) in C3 images, and quite diffuse with a very faint short tail in C2 images. C/2008 R12 was slightly elongated (mag ∼ 6–6.5) in C3 images, and diffuse with a very faint, diffuse, short tail in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 R8 Sept. 3.629 10^{h}46.^{m}2 + 6^{o}02′ C2  RK 2008-X24
C/2008 R9 4.904 10 32.1 + 5 34 C3 JR 2008-X24
C/2008 R10  7.904 10 38.0 + 4 00 C3/2 RK 2008-X24
C/2008 R11 7.904 10 43.1 + 4 35 C3 BZ 2008-X25
C/2008 R12  11.905 10 54.5 + 2 40 C3/2 BZ 2008-X25
C/2008 R13  14.596 11 16.4 + 2 29 C3  HS 2008-X25
2009 January 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green